# Predictive Linear Online Tracking for Unknown Targets

### v1

@@ 2.2 LQT Optimal Controller
Having full access to future target states, the optimal controller admits the following analytical solution Foster and Simchowitz (2020); Goel and Hassibi (2022)
@@
$K$ $=(R+B^{\top}PB)^{-1}B^{\top}PA,$ (8)
$K_{t}$ $=(R+B^{\top}PB)^{-1}B^{\top}(A-BK)^{\top,t}P,t\geq 0,$ (9)
with $P$ defined in (6). Since the dynamics (1) are known, all gain matrices $K,K_{t},\,t\geq 0$ are known. The feedback term $-K(x_{t}-r_{t})$ can be computed since the error $x_{t}-r_{t}$ is measured before choosing the action. The only unknown is the feedforward term $q_{t}(r_{t:T})$. Note that the functional form of the feedforward term is known. The missing piece of information is the future target states. Following the paradigm of indirect control, we will use this structural observation to decompose the problem of online tracking into online prediction and control design.
@@
Upper bounds for $c_{0},\rho$ as well as a relaxed version of Assumption 2.3 can be found in Appendix A. Interestingly, the above property implies that future target states get discounted when considering the current action; we only need to know accurately the imminent target states.
@@ 3 Predictive Linear Online Tracking
Target Prediction
For the prediction step, we employ the RLS algorithm with forgetting factors. At every time step $t$, the RLS algorithm provides predictions $r_{t+1|t},\dots,r_{t+W|t}$ of the future target states $r_{t+1},\dots,r_{t+W}$, for some horizon $W>0$. When predicting the $k-$step ahead target $r_{t+k|t}$, we only receive the true value $r_{t+k}$ after $k-$steps into the future. To deal with this “delayed feedback” issue we follow the approach of Joulani et al. (2013). The estimates $\hat{S}_{t+k|t}$ of $S_{t+k|t}$ are handled by $k$ independent learners that are updated at non-overlapping time steps. For example, the estimate $\hat{S}_{t+k|t}$ is updated based on pairs of $(r_{j},z_{j-k})$ collected at times $j=t,t-k,\dots$. At each time step $t$, only one of the $k$ independent learners is invoked, which aims to minimize the following prediction error $\forall t>1$:
@@
where $\gamma\in(0,1]$ is the forgetting factor that attributes higher weights to more recent data points. Recall that $\mathcal{S}$ is the set in which the dynamics lie–see Assumption 2.2. Note that each learner uses at most $\lceil T/k\rceil$ data points. Instead of solving (11), we opt for a recursive implementation with projections adapted from Yuan and Lamperski (2020) and can be found in Algorithm 1.
@@
0:  Forgetting factor $\gamma\in(0,1)$, Regularizer $\epsilon$
@@
Different from direct approaches Foster and Simchowitz (2020); Baby and Wang (2022), where the learning feedback is delayed by $W$ steps, here the learning feedback delay adapts to the $k-$step ahead horizon offering more flexibility. For example, the feedback for the one-step ahead predictor is always available without waiting $W$ steps. A downside is that we update more learners. Note that we follow an improper learning approach. Instead of learning over matrices $S_{t+W},\dots,S_{t+1}$, we learn over the space of multi-step ahead predictors $S_{t+W|t},\dots,S_{t+1|t}$ ignoring their inter-dependence. The former parameterization leads to non-convex optimization problems since the multi-step ahead predictors are products of entries in $S_{t+W},\dots,S_{t+1}$. The latter parameterization leads to a higher-dimensional but convex problem.
Receding Horizon Control
In the receding horizon control framework, instead of looking over the full control horizon $T$ as in (2), we only look over a window of length $W$ and solve the following optimal control problem
@@
1:  Compute $P,K,\{K_{0},...,K_{W-1}\}$ as in (6), (8), (9).
2:  Initialize RLS learners according to the Algorithm 1, for $k=1,...,W$, respectively.
@@
7:           Update the $k$-step-ahead learner and predict $r_{t+k|t}$ according to the Algorithm 1.
@@
The main part of our proposed algorithm, Predictive Linear Online Tracking (PLOT), is described in Algorithm 2. We initialize $W$ instances of the RLS (Algorithm 1) to estimate the $k$-step-ahead dynamics $S_{t+k|t}$ for $k=1,...,W$ independently. Since we need $k$ learners for every $k-$step ahead predictor, we have a total number of $W(W+1)/2$ learners. At each time step, the algorithm receives the latest target and state measurements, it updates the RLS learners, and makes new predictions. Then, it computes the receding horizon control input given the predictions and the state information. The whole procedure is iterated over all time steps up to the control horizon $T$. Note that if the prediction time $k+t$ exceeds the control horizon $T$, then we set the predicted target to zero; as seen by (7) only the target states up to time $T$ affect the control problem.
@@ 4 Dynamic Regret and Tuning
$-\alpha_{3}\tilde{W}^{2}(T+1)\log\gamma-\alpha_{4}\tilde{W}^{3}\log(1-\gamma)+\alpha_{5}\tilde{W}^{3},$
@@
The first term in the regret bound captures the truncation effect, i.e., the fact that we only use $W$-step (instead of $T$) ahead predictions in (12). The other terms capture the effect of prediction errors on the control performance. Note that there is a tradeoff between large and small prediction horizons. Smaller prediction horizons lead to better prediction performance but worse truncation error and, conversely, larger prediction horizons improve the truncation term but degrade the prediction performance. Nonetheless, thanks to the exponentially decaying LQT gains, if we increase the prediction horizon $W$ past the threshold of $(1-\rho)^{-1}$, then the regret guarantees stop degrading. This is a consequence of using multiple predictors of varying delay and improves prior work Foster and Simchowitz (2020); Baby and Wang (2022). Essentially, $\tilde{W}$ can be thought of as the “effective” prediction horizon; beyond $\tilde{W}$ any prediction errors have negligible effect.
Assume that $\gamma$ is close to $1$. Then $-\log\gamma\approx 1-\gamma$ and the dominant regret terms are the second and third terms, $V(1-\gamma)^{-1}$ and $T(1-\gamma)$ respectively. By balancing these two terms, we obtain the following interpretable rate, similar to Yuan and Lamperski (2020).
@@
To deal with the truncation term, it is sufficient to choose a prediction horizon $W$, which grows logarithmically with the control horizon $T$, as also noted in Zhang et al. (2021b). The constant $4M$ in the denominator guarantees that the forgetting factor is positive. To tune the forgetting factor and obtain the rate, we require knowledge of the path length $V_{T}$, which is not always available. Following the procedure of Baby and Wang (2022), we can overcome this limitation by initiating multiple learners at various time steps and by running a follow-the-leading-history (FLH) meta-algorithm on top, which can also improve the dynamic regret bound to $T^{1/3}V_{T}^{2/3}$. We do not explore this possibility here, since we focus on studying the performance of the forgetting factor RLS algorithm. When the path length is close to 0, i.e. the target dynamics are almost static, we obtain logarithmic regret guarantees. In this case, the optimal policy can be rewritten as a static affine control law, with respect to past target states. The regret, in this case, is the static regret with respect to the best (static) affine policy, recovering the result of Foster and Simchowitz (2020).
@@
In certain works Li et al. (2019); Karapetyan et al. (2023), the dynamic regret is given with respect to the total variation of the target states themselves, that is, $L_{T}=\sum_{t=2}^{T-1}\lVert r_{t}-r_{t-1}\rVert$. In contrast, here, we capture learning complexity by the total variation of the target dynamics $V_{T}$. We argue that the former notion of complexity can be suboptimal in the case of targets with dynamic structure, while ours is superior in this setting. For example, consider a target that is moving on a circle with constant velocity. Then, the former path length is linear $L_{T}=\mathcal{O}(T)$ while the latter is zero $V_{T}=0$. More discussion can be found in Appendix B.1.
The full proof of Theorem 4.1 and Corollary 4.2 can be found in the Appendix. In the following, we provide a sketch of the proof. First, similar to Foster and Simchowitz (2020), we invoke the “performance difference lemma” Kakade (2003) to turn the control problem into a prediction problem.  [^4] In our case, the optimal controller is the true minimizer of the cost, hence, the regret in this case becomes the advantage function of Foster and Simchowitz (2020).
Lemma 4.4 (Performance Difference Lemma Foster and Simchowitz (2020)).
@@
$\mathcal{R}(\pi)=\sum_{t=0}^{T-1}\lVert\hat{q}_{t}-q_{t}(r_{t:W})\rVert^{2}_{(R+B^{T}PB)}.$
The above fundamental result shows that the control problem can be cast as a prediction problem, where we try to predict the affine part of the optimal control law. Second, we bound the prediction error by analyzing the dynamic regret of the RLS algorithm following the steps of Yuan and Lamperski (2020). Define the prediction regret for $k-$step ahead prediction as
@@
To conclude the proof, we combine the previous steps while also accounting for the truncation effect. The only remaining step is to upper bound the $k$-steps ahead path lengths $V^{k}_{T}$ in terms of the path length of the one-step-ahead path length $V_{T}$ as defined in (14). In Lemma B.3, we prove that $V^{k}_{T}\leq\sqrt{p}k^{2}M^{2}V_{T}$. This is why the factor $\tilde{W}^{4}$ that multiplies the path length in the final regret bound in Theorem 4.1 appears with an exponent of $4$ instead of just $2$.
@@ 5 Simulations and Experimental Validation
In this section, we demonstrate PLOT’s performance in simulation and hardware experiments and provide the corresponding code in https://gitlab.nccr-automation.ch/akarapetyan/plot. We study the setting of tracking adversarial unknown targets with a quadrotor, which is of particular interest given the number of potential applications, such as hunting adversarial drones in airports Dressel and Kochenderfer (2019) or artistic choreography Schoellig et al. (2014).
@@
We consider the Crazyflie 2.1 quadrotor Bitcraze (2023b), a versatile, open-source, nano-sized quadrotor developed by Bitcraze Bitcraze (2023b), whose dynamics can be modeled with a linear time-invariant system linearized around a hovering point Beuchat (2019). For such a model, we define the state ${x:=[\vec{p}~{};~{}\vec{\dot{p}}~{};~{}\vec{\psi}]}$, where ${\vec{p}:=\begin{bmatrix}x&y&z\end{bmatrix}^{\top}}$ is the position vector in the inertial frame, ${\vec{\psi}:=\begin{bmatrix}\gamma&\beta&\alpha\end{bmatrix}^{\top}}$ is the attitude vector in the inertial frame with $\gamma$, $\beta$ and $\alpha$ for the roll, pitch and yaw angle, respectively. The action ${u:=[f~{};~{}\vec{\omega}]}$ includes the total thrust $f$ and the angular rate ${\vec{\omega}:=\begin{bmatrix}\omega_{x}&\omega_{y}&\omega_{z}\end{bmatrix}^{\top}}$ in the body frame. For the detailed derivation of the linear dynamic model, see Appendix E and Beuchat (2019). For all examples, we take the sampling time to be $T_{s}=0.1$ seconds and the LQR cost matrices fixed at $Q:=diag(80,80,80,10,10,10,0.01,0.01,0.1)$, and $R:=diag(0.7,2.5,2.5,2.5)$.
@@ 5.1 Simulation Results
Here, we consider the linearized dynamics of the Crazyflie quadrotor, derived in Appendix E.1. We show how PLOT tracks a circular reference target with $V_{T}=0$ path length per (14), we also provide performance comparisons to other online control algorithms for tracking a much harder reference with $V_{T}=\sqrt{T}$.

### v3

@@ 2.2 LQT Optimal Controller
Having full access to future target states, the optimal controller admits the following analytical solution (Foster & Simchowitz, 2020; Goel & Hassibi, 2022)
@@
$K$ $=(R+B^{\top}XB)^{-1}B^{\top}XA,$ (8)
$K_{t}$ $=(R+B^{\top}XB)^{-1}B^{\top}(A-BK)^{\top,t}X,\;t\geq 0,$ (9)
with $X$ defined in (6). Since the dynamics (1) are known, all gain matrices $K,K_{t},\,t\geq 0$ are known. The feedback term $-K(x_{t}-r_{t})$ can be computed since the error $x_{t}-r_{t}$ is measured before choosing the action. The only unknown is the feedforward term $q_{t}(r_{t:T})$. Note that the functional form of the feedforward term is known. The missing piece of information is the future target states. Following the paradigm of indirect control, we will use this structural observation to decompose the problem of online tracking into one of online prediction and control design.
@@
Upper bounds for $c_{0},\rho$ as well as a relaxed version of Assumption 2.3 can be found in Appendix B. Interestingly, the above property implies that future target states get discounted when considering the current action; we only need to know accurately the imminent target states.
@@ 3 Predictive Linear Online Tracking
Target Prediction.
For the prediction step, we employ the RLS algorithm with forgetting factors. At every time step $t$, the RLS algorithm provides predictions $r_{t+1|t},\dots,r_{t+W|t}$ of the future target states $r_{t+1},\dots,r_{t+W}$, for some horizon $W>0$. For every step ahead prediction, we keep $W$ separate predictors $\hat{S}_{t+1|t}$,…, $\hat{S}_{t+W|t}$ without exploiting the shared structure. This parameterization leads to a higher-dimensional but convex problem. In contrast, learning over matrices $S_{t+W},\dots,S_{t+1}$ leads to non-convex optimization problems since the multi-step ahead predictors are products of entries in $S_{t+W},\dots,S_{t+1}$–see also Remark 3.1.
When predicting the target state $r_{t+k|t}$ for $k$ steps ahead, the true value $r_{t+k}$ is revealed after $k$ steps into the future. To deal with this delayed feedback issue we follow the approach of Joulani et al. (2013). Within each $k-$step ahead predictor, we maintain $k$ copies $\{\hat{S}_{t+k-j|t-j}\}^{k-1}_{j=0}$–we refer to them as learners–which are updated independently at non-overlapping time steps. Hence, the estimate $\hat{S}_{t+k|t}$ is updated based only on pairs of $(r_{t-ik},z_{t-(i+1)k})$, for $i=0,1,2,\dots$. At each time step $t$, only one of the $k$ independent learners is invoked, which aims to minimize the following prediction error $\forall t>1$:
@@
where $\gamma\in(0,1]$ is the forgetting factor that attributes higher weights to more recent data points. Recall that $\mathcal{S}$ is the set in which the dynamics lie–see Assumption 2.2. Note that each independent learner is updated based on at most $\lceil T/k\rceil$ data points. Instead of solving (11), we opt for a recursive implementation with projections adapted from Yuan & Lamperski (2020) and can be found in Algorithm 1.
@@
0:  Forgetting factor $\gamma\in(0,1)$, Regularizer $\varepsilon>0$.
@@
Since we need $k$ learners for every $k-$step ahead predictor, we have a total number of $W(W+1)/2$ learners. However, only $W$ of them are actively updated at every time step (one active learner per predictor). An visualization of the learning architecture for $k=1,2,3$ can be found in Figure 7 in Section D.2.
Different from direct approaches (Foster & Simchowitz, 2020; Baby & Wang, 2022), where the learning feedback is delayed by $W$ steps, here the learning feedback delay adapts to the $k-$step ahead horizon offering more flexibility. For example, the feedback for the one-step ahead predictor is always available without waiting $W$ steps. A downside is that we update more predictors.
Remark 3.1 (Improper Learning versus Single Learner).
To obtain the multi-step ahead predictions $\hat{r}_{t+k|t}$, $k=1,\dots,W$, we keep $W$ separate predictors $\hat{S}_{t+1|t}$,…, $\hat{S}_{t+W|t}$ without exploiting the shared structure, i.e., we follow improper learning. By following this choice, we convexify the problem, potentially at the expense of higher sample complexity. An alternative, “naive” approach is to estimate the one-step ahead predictor $\hat{S}_{t+1|t}$, using a single learner. Then, to predict $k$ steps ahead, e.g. $\hat{r}_{t+k|t}$, we could propagate the target states through $\hat{S}_{t+1}$, for $k$ time steps. However, propagating through multiple time steps can lead to instability. For example, if $z_{t}=r_{t}$ (namely the AR memory length $p=1$), we will have $\hat{r}_{t+k|t}=\hat{S}^{k}_{t+1}r_{t}$. Due to the exponent $k$, this could lead to a rapid accumulation of errors. Enforcing the boundedness constraint of Assumption 2.2 directly on $\hat{S}^{k}_{t+1}$ would not work either since it leads to a non-convex problem. A detailed comparison of PLOT and this “naive” approach in simulation can be found in Appendix F.4
Receding Horizon Control.
In the receding horizon control framework, instead of looking over the total time $T$ as in (2), we only look over a window of length $W$ and solve the following optimal control problem
@@
1:  Compute $X,K,\{K_{0},...,K_{W-1}\}$ as in (6), (8), (9).
2:  Initialize RLS predictors according to the Algorithm 1, for $k=1,...,W$, respectively.
@@
7:           Update the $k$-step-ahead predictor and predict $r_{t+k|t}$ according to the Algorithm 1.
@@
The main part of our proposed algorithm, Predictive Linear Online Tracking (PLOT), is described in Algorithm 2. We initialize $W$ RLS predictors (Algorithm 1) to estimate the $k$-step-ahead dynamics $S_{t+k|t}$ for $k=1,...,W$ independently. At each time step, the algorithm receives the latest target and state measurements, updates the RLS predictors, and makes new predictions. Then, it computes the receding horizon control input given the predictions and the state information. The whole procedure is iterated over all time steps up to the total time $T$. Note that if the prediction time $k+t$ exceeds the total time $T$, then we set the predicted target to zero; as seen by (7) only the target states up to time $T$ affect the control problem.
@@ 4 Dynamic Regret and Tuning
$\quad-\alpha_{3}\tilde{W}^{2}(T+1)\log\gamma-\alpha_{4}\tilde{W}^{3}\log(1-\gamma)+\alpha_{5}\tilde{W}^{3},$
@@
The first term in the regret bound captures the truncation effect, i.e., the fact that we only use $W$-step (instead of $T$) ahead predictions in (12). The other terms capture the effect of prediction errors on the control performance. Note that there is a tradeoff between large and small prediction horizons. Smaller prediction horizons lead to better prediction performance but worse truncation error and, conversely, larger prediction horizons improve the truncation term but degrade the prediction performance. Nonetheless, if we increase the prediction horizon $W$ past the threshold of $(1-\rho)^{-1}$, then the regret guarantees stop degrading, thus improving prior work (Foster & Simchowitz, 2020; Baby & Wang, 2022). Essentially, $\tilde{W}$ can be thought of as the “effective” prediction horizon; beyond $\tilde{W}$ any prediction errors have negligible effect. This is a consequence of deploying multiple step-ahead predictors of varying delays and the exponentially decaying LQT gains,
Assume that $\gamma$ is close to $1$. Then $-\log\gamma\approx 1-\gamma$ and the dominant regret terms are the second and third terms, $V(1-\gamma)^{-1}$ and $T(1-\gamma)$ respectively. By balancing these two terms, we obtain the following interpretable rate, similar to Yuan & Lamperski (2020).
@@
To deal with the truncation term, it is sufficient to choose a prediction horizon $W$, which grows logarithmically with the total time $T$, as also noted in Zhang et al. (2021b). The constant $4M$ in the denominator guarantees that the forgetting factor is positive. To tune the forgetting factor and obtain the rate, we require knowledge of the path length $V_{T}$, which is not always available. Following the procedure of Baby & Wang (2022), we can overcome this limitation by initiating multiple learners at various time steps and by running a follow-the-leading-history (FLH) meta-algorithm on top, which can also improve the dynamic regret bound to $T^{1/3}V_{T}^{2/3}$. We do not explore this possibility here, since we focus on studying the performance of the RLS algorithm with forgetting factors. When the path length is close to 0, i.e. the target dynamics are almost static, we obtain logarithmic regret guarantees. In this case, the optimal policy can be rewritten as a static affine control law, with respect to the past target states. The regret, in this case, is the static regret with respect to the best (static) affine policy, recovering the result of Foster & Simchowitz (2020).
@@
In certain works (Li et al., 2019; Karapetyan et al., 2023), the dynamic regret is given with respect to the total variation of the target states themselves, that is, $L_{T}=\sum_{t=2}^{T-1}\lVert r_{t}-r_{t-1}\rVert$. In contrast, here, we capture learning complexity by the total variation of the target dynamics $V_{T}$. We argue that the former notion of complexity can be suboptimal in the case of targets with dynamic structure, while ours is superior in this setting. For example, consider a target that is moving on a circle with constant velocity–see C.2. Then, the former path length is linear $L_{T}=\mathcal{O}(T)$ while the latter is zero $V_{T}=0$. In control applications, many targets of interest have dynamic structures, e.g. tracking other drones, tracking moving objects, etc. Nonetheless, if the targets are fully unstructured and arbitrary, PLOT might not perform as well since it tries to fit a dynamic model to arbitrary data–see Appendix F.3.2. Still, the regret order cannot be worse than the one in prior work if we are learning over ARX models–see discussion in Appendix C.1.
The full proof of Theorem 4.1 and Corollary 4.2 can be found in Appendix E. In the following, we provide a sketch of the proof. First, similar to Foster & Simchowitz (2020), we invoke the “performance difference lemma” (Kakade, 2003) to turn the control problem into a prediction one  [^1] In our case, the optimal controller is the true minimizer of the cost, hence, the regret in this case becomes the advantage function of (Foster & Simchowitz, 2020)..
Lemma 4.4 (Performance Difference Lemma (Foster & Simchowitz, 2020)).
@@
$\mathcal{R}(\pi)=\sum_{t=0}^{T-1}\lVert\hat{q}_{t}-q_{t}(r_{t:W})\rVert^{2}_{(R+B^{T}XB)}.$
The above fundamental result shows that the control problem can be cast as a prediction problem, where we try to predict the affine part of the optimal control law. Second, we bound the prediction error by analyzing the dynamic regret of the RLS algorithm following the steps of Yuan & Lamperski (2020). Define the prediction regret for $k-$step ahead prediction as
@@
To conclude the proof, we combine the previous steps while also accounting for the truncation effect. The only remaining step is to upper bound the $k$-steps ahead path lengths $V^{k}_{T}$ in terms of the path length of the one-step-ahead path length $V_{T}$ as defined in (14). In Lemma C.3, we prove that $V^{k}_{T}\leq\sqrt{p}k^{2}M^{2}V_{T}$. This is why the factor $\tilde{W}^{4}$ that multiplies the path length in the final regret bound in Theorem 4.1 appears with an exponent of $4$ instead of just $2$.
@@ 5 Simulations and Experimental Validation
In this section, we demonstrate the performance of PLOT both in simulation and in hardware experiments. We study the setting of tracking adversarial unknown targets with a quadrotor, which is of particular interest given the number of potential applications, such as hunting adversarial drones in airports (Dressel & Kochenderfer, 2019) or artistic choreography (Schoellig et al., 2014), to name a few. The code for both the simulation and hardware experiments can be accessed from https://gitlab.nccr-automation.ch/akarapetyan/plot.
@@
We consider the Crazyflie 2.1 quadrotor (Bitcraze, 2023b), a versatile, open-source, nano-sized quadrotor developed by Bitcraze (Bitcraze, 2023b), whose dynamics can be modeled with a linear time-invariant system linearized around a hovering point (Beuchat, 2019). For such a model, we define the state ${x:=[\vec{p}~{};~{}\vec{\dot{p}}~{};~{}\vec{\psi}]}$, where ${\vec{p}:=\begin{bmatrix}x&y&z\end{bmatrix}^{\top}}$ is the position vector in the inertial frame, ${\vec{\psi}:=\begin{bmatrix}\gamma&\beta&\alpha\end{bmatrix}^{\top}}$ is the attitude vector in the inertial frame with $\gamma$, $\beta$ and $\alpha$ for the roll, pitch and yaw angle, respectively. The action ${u:=[f~{};~{}\vec{\omega}]}$ includes the total thrust $f$ and the angular rate ${\vec{\omega}:=\begin{bmatrix}\omega_{x}&\omega_{y}&\omega_{z}\end{bmatrix}^{\top}}$ in the body frame. For the detailed derivation of the linear dynamic model, see Appendix F.1 and (Beuchat, 2019). For all examples, we take the sampling time to be $T_{s}=0.1$ seconds and the LQR cost matrices fixed at $Q:=diag(80,80,80,10,10,10,0.01,0.01,0.1)$, and $R:=diag(0.7,2.5,2.5,2.5)$.
@@ 5.1 Simulation Results
Here, we consider the linearized dynamics of the Crazyflie quadrotor, derived in Appendix F.1.We show how PLOT successfully tracks both a static reference target with $V_{T}=0$ path length and a dynamic target with $V_{T}=\mathcal{O}(\sqrt{T})$. In addition, we also demonstrate how the proven regret guarantees can be used as a guideline for tuning the algorithm parameters such as the prediction horizon length $W$ and the forgetting factor $\gamma$. We also provide a performance comparison to other online control algorithms from the literature for tracking a non-stationary target.